Circular No. 8811

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(123509) 2000 WK_{183}

K. S. Noll, Space Telescope Science Institute (STScI); D. C. Stephens, Brigham Young University; W. M. Grundy, Lowell Observatory; H. F. Levison, Southwest Research Institute; and S. D. Kern, STScI, report the detection of a binary companion to the transneptunian object (123509) in observations made during 2005 Nov. 24.365–24.395 UT with the High Resolution Camera of the Advanced Camera for Surveys on the Hubble Space Telescope (HST), using the clear filters with one 300-s exposure at each of four dithered positions on the detector. The two components are clearly resolved in each image and in the coadded image; they were separated by an angular distance of 0′′.080 ± 0′′.004 and differ in brightness by 0.40 magnitude. The fainter component lies at a position angle of 138^{o}.9 ± 5^{o}.5 from the primary. The projected separation of the objects in the sky plane is 2470 ± 60 km. The HST corrected for parallax and tracked both components of (123509) as they moved together at an average rate of 0′′.50/min.

COMETS C/2006 V2–V10 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8807). C/2006 V2 was very bright, peaking at mag 1.3 on Nov. 3.154 UT at 11.0R_{⊙} in C3 images, with a thin tail that was 1^{o}.25 long at 8.0R_{⊙} on Nov. 3.429; in C2 images, the long, thin tail survived for some hours after the comet’s head had disappeared behind the occulter. C/2006 V3 and V4 were stellar in appearance and faint (mag 7.5). C/2006 V5, V6, V7, V8, and V10 were somewhat diffuse and faint (mag ∼ 8, 7.5, 7, 7.5, and 8, respectively). C/2006 V9 was stellar in appearance and of mag ∼ 5 in C3 images; in C2 images, it had a condensed head and an extremely faint, thin tail that extended to ∼ 17′ on Nov. 14.171 at 5R_{⊙}.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 V2 Nov. 2.013 14^{h}00.^{m}6 −18^{o}18′ C3/2 JS 2007-B38
C/2006 V3  3.021 14 27.2 −16 13 C2  HS 2007-B38
C/2006 V4  4.013 14 25.3 −17 28 C3  HS 2007-B38
C/2006 V5  7.379 14 46.0 −17 44 C2  HS 2007-B38
C/2006 V6  7.521 14 46.1 −17 43 C2  TH 2007-B38
C/2006 V7  8.579 14 50.6 −18 02 C2  HS 2007-B38
C/2006 V8  11.851 15 04.7 −19 07 C2  RK 2007-B75
C/2006 V9  13.321 14 59.3 −21 16 C3/2 BZ 2007-B75
C/2006 V10  13.993 15 13.4 −19 40 C2  HS 2007-B75
2007 February 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green